OpenStudy (anonymous):

Solve for y (4y+5/3)=(y-10/2)

OpenStudy (anonymous):

\[\frac{4y+5}{3}=\frac{y-10}{2}\]

OpenStudy (anonymous):

cross and time

OpenStudy (anonymous):

2(4y+5)=3(y-10) 8y+10=3y-30 8y-3y= -30-10 5y = -40 y = -40/5 y = -8